# Small-Scale and Transient EUV Kernels in Solar Flare Ribbons

### 12 Mar 2026 v1

@@ 4.1 Kernel areas
HRIwas designed for pixel-limited resolution (Rochus, P. et al., 2020) and early observations (Fig. 24, Berghmans et al., 2023) indeed indicate that the sharpest observable features are close to the Nyquist Sampling limit of 2-pixels. In Appendix A, Figure 9, an example is given that shows a point source smeared out by a PSF with a 2-pixel FWHM. The point source is centred in the middle of a pixel and at the intersection between four pixels in the top and bottom row, respectively. If the thresholds are not chosen well, the method applied in this work can lead to kernel shrinkage. This effect is demonstrated in the final three columns of Figure 9 where three cut-off thresholds were applied to a single kernel. It is clear that a point source with this PSF may appear to have anywhere between 1–20 pixels depending on the position of the centre of the source on the pixel and the cut-off threshold applied. Since the threshold is a fixed value, fainter kernels are more susceptible to artificial source shrinkage.
@@
Figure 5 indicates that a significant fraction of EUV kernels were unresolved when observed with the instrumental resolution offered by HRIat this radial distance. To quantify this, we excluded kernels that contain less than 10 pixels from our analysis because we know that those are affected by artificial shrinkage. This amounted to 50% of the total kernels when the 10% threshold was applied to the original data. We estimated that structures with 10–20 pixels correspond to roughly 1 PSF (see Appendix A). Of the remaining 50%, 50% contained 10–20 pixels when segmented on the original data using a threshold of 10%. The fraction decreased to 43% of kernels when a 5% threshold was applied. Using the “running difference” data, the percentages increased to 73% and 60% when the 10% and 5% thresholds were applied, respectively. This highlights the fact that energy deposition occurred on scales unresolved by the instrument in a significant fraction of events. Figure 5 a) also shows the kernel size distribution of the 20% brightest kernels (based on their maximum intensity). The mean of the distribution is greater than that obtained when the entire range of kernel intensities was included and there is a turnover in the distribution at small kernel sizes. This is partly attributable to the effect of kernel shrinkage due to the application of thresholds to the data. However, Figure 6 also shows that there is also a strong positive correlation between the peak kernel intensity and the size of the kernel. Excluding kernels that consisted of less than 10 pixels, Spearman and Pearson correlation coefficients, and were obtained, with p-values, , in both cases. This indicates that there is a fundamental scaling between the energetics of the injected flux and the spatial extent of heating in the solar atmosphere. Based on the data obtained, we cannot definitively determine how this scaling relation extends to the lower range, but if we extrapolate the existing trend, there would be even more small, faint kernels, that would increase the estimated fraction of those that were unresolved.
@@
Figure 7 a) shows the total short-exposure EUV intensity (in ) versus the total number of bright pixels in each frame, identified by the watershed algorithm when the 5% threshold was applied. There is a strong positive correlation (, for both) between the total intensity of the EUV ribbons (in ) and the total number of kernel pixels in a frame. In other words, the total instantaneous EUV ribbon intensity (shown in panel b)) correlates with the instantaneous ribbon area (shown in panel c)). This result is surprising because, if the relationship were perfectly linear, it would indicate that the EUV flux was evenly distributed across kernels in time. However, panel d) shows the total DN/s normalised by the instantaneous ribbon area. Interestingly, the time evolution corresponds precisely to the time evolution of the spatially integrated EUV ribbon intensity (panel b)). In other words, the temporal evolution is still present in the EUV flux. As a result, we conclude that the total ribbon intensity correlates with the ribbon area to a certain degree; however, there was also variation in the energy flux within certain regions, that led to the temporal evolution shown in panel d). This finding is largely consistent with Figure 6 which examines the relationship between the peak intensity of individual kernels versus their size.
@@ 6 Conclusions
In this work, data from the 2024 major flare campaign of Solar Orbiter were analysed. Those data showed that during the impulsive phase of the M2.5 GOES-class flare, SOL2024-03-23T23:41, when electron acceleration was strong, a significant fraction (approximately 50%) of the EUV kernels measured by HRIin the 174 Å passband were unresolved at a pixel scale of 135 km/pix and 2-pixel FWHM of the PSF core. These results are consistent with recent observations of chromospheric blobs in flare ribbons at high spatial resolution with SST and the DKIST (Thoen Faber et al., 2025; Yadav et al., 2025). In those works, chromospheric blobs with FWHM sizes in the range 140–200 km and 320–455 km were identified by SST and the DKIST, respectively. This corresponds to between 1–4 HRIpixels at 0.38 AU. Structures of this size are not resolvable by HRIEUV. Our results therefore agree with those obtained by SST and the DKIST, depsite the fact that HRIprobes a different layer of the solar atmosphere than the SST and the DKIST. In addition, we showed that for the flare studied, energy deposition occurred in a new region along the flare arcade in quick succession (on timescales of several seconds). A strong positive correlation was found between the peak kernel intensity and kernel size, highlighting a fundamental relationship between the energetics of flare energy release and the spatial extent of the event. Furthermore, an average EUV kernel time profile was derived. An upper bound was placed on the average half-maximum heating time of only a couple of seconds in duration, with correspondingly fast cooling times. Our best fit of the average EUV kernel time profile gave and . Together, these results highlight the transient and fragmented nature of solar flare energy release on small spatial scales. The findings presented here have important implications on our understanding of flare energy deposition. The full extent to which these observations alter our understanding warrants detailed modelling investigation. However, what remains strikingly clear is that the energy release process in flares is highly fragmented in space and time. Finally, we conclude that the bulk energy release process of solar flares consists of many smaller scale () events that behave like multiple flare-like energy release episodes occurring in quick succession.
@@ Acknowledgements.
We thank the anonymous referee for their constructive feedback and insightful comments. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. The EUI instrument was built by CSL, IAS, MPS, MSSL/UCL, PMOD/WRC, ROB, LCF/IO with funding from the Belgian Federal Science Policy Office (BELSPO/PRODEX PEA 4000112292 and 4000134088); the Centre National d’Etudes Spatiales (CNES); the UK Space Agency (UKSA); the Bundesministerium für Wirtschaft und Energie (BMWi) through the Deutsches Zentrum für Luft- und Raumfahrt (DLR); and the Swiss Space Office (SSO). HC and LAH had additional support from an EUI guest investigator grant issued by the EUI instrument team. The STIX instrument is an international collaboration between Switzerland, Poland, France, Czech Republic, Germany, Austria, Ireland, and Italy. LAH is supported by a Royal Society-Research Ireland University Research Fellowship (URFR1241775).References
@@
Appendix A A point source observed with a 2-pixel PSF
In Figure 9 we demonstrate the effect of a 2-pixel PSF on a point source located at two extreme positions — at the centre of a pixel and at the intersection of four pixels (top and bottom row, respectively). The effect of applying a threshold is demonstrated in the final three columns. Naturally, as the threshold increases the number of pixels with non-zero counts reduces, decreasing the size of the kernel that would be identified from processed data. In this work, a fixed cut-off threshold is used for the entire time series. As a result, fainter kernels are more affected by artificial shrinkage. We therefore caution the reader not to over-interpret the shape of the size distributions shown in Figure 5 in the low-intensity range. This highlights the fact that point sources that have been artificially shrunk due to data preprocessing may appear as kernels with up to 9 pixels. Further, kernels with between 10 and 20 pixels are likely unresolved by HRIEUV.
Appendix B Average EUV kernel time profiles

### 02 Apr 2026 v2

@@ 4.1 Kernel areas
HRIwas designed for pixel-limited resolution (Rochus, P. et al., 2020) and early observations (Fig. 24, Berghmans et al., 2023) indeed indicate that the sharpest observable features are close to the Nyquist Sampling limit of 2-pixels. In Appendix A, Figure 8, an example is given showing a point source smeared out by a PSF with a 2-pixel FWHM. The point source is centred in the middle of a pixel and at the intersection between four pixels in the top and bottom row, respectively. If the thresholds are not chosen well, the method applied in this work can lead to kernel shrinkage. This effect is demonstrated in the final four columns of Figure 8 where four cut-off thresholds were applied to a single kernel. It is clear that a point source with this PSF may appear to have anywhere between 1–20 pixels depending on the position of the centre of the source on the pixel and the cut-off threshold applied. Since the threshold is a fixed value, fainter kernels are more susceptible to artificial source shrinkage.
@@
Figure 5 indicates that a significant fraction of EUV kernels were unresolved when observed with the instrumental resolution offered by HRIat this radial distance. To quantify this, we excluded kernels that contain less than 10 pixels from our analysis because we know that those are affected by artificial shrinkage. This amounted to 50% of the total kernels when the 10% threshold was applied to the original data. We estimated that structures with 10–20 pixels correspond to roughly 1 PSF (see Appendix A). Of the remaining 50%, 50% contained 10–20 pixels when segmented on the original data using a threshold of 10%. The fraction decreased to 43% of kernels when a 5% threshold was applied. Using the “running difference” data, the percentages increased to 73% and 60% when the 10% and 5% thresholds were applied, respectively. This highlights the fact that energy deposition occurred on scales unresolved by the instrument in a significant fraction of events. Figure 5 a) also shows the kernel size distribution of the 20% brightest kernels (based on their maximum intensity). The mean of the distribution is greater than that obtained when the entire range of kernel intensities was included and there is a turnover in the distribution at small kernel sizes. This difference is partly attributable to the effect of kernel shrinkage due to the application of thresholds to the data. In Appendix B we also show the size distributions for a subset of kernels whose watershed labels extend beyond pixels with intensities equal to 50% of the peak kernel intensity. The labels were clipped to only include pixels with intensities . This removes any differences introduced by using a fixed threshold. See Figure 9 for a comparison with Figure 5. Since the kernels identified in this dataset, are close to and sometimes at the resolution limit of the instrument, the source size estimates are even sensitive to where the source falls on the pixel grid, as demonstrated in Figure 8. Kernels are not smoothly sampled and have steep gradients. As a result, it is challenging to determine the true source sizes and we therefore caution the reader not to over-interpret the shape of the distributions presented in Figure 5 and Figure 9. The figures primarily serve to highlight the main range of kernel sizes detected by HRIin this dataset. We also note that a very small fraction () of kernels detected contained more than pixels (not shown in Figure 5).
@@
Figure 6 a) shows the total short-exposure EUV intensity (in ) versus the total number of bright pixels in each frame, identified by the watershed algorithm when the 5% threshold was applied. There is a strong positive correlation (, for both) between the total intensity of the EUV ribbons (in ) and the total number of kernel pixels in a frame. In other words, the total instantaneous EUV ribbon intensity (shown in panel b)) correlates with the instantaneous ribbon area (shown in panel c)). This result is surprising because, if the relationship were perfectly linear, it would indicate that the EUV flux was evenly distributed across kernels in time. However, panel d) shows the total DN/s normalised by the instantaneous ribbon area. Interestingly, the time evolution corresponds precisely to the time evolution of the spatially integrated EUV ribbon intensity (panel b)). In other words, the temporal evolution is still present in the EUV flux. As a result, we conclude that the total ribbon intensity correlates with the ribbon area to a certain degree; however, there was also variation in the energy flux within certain regions, that led to the temporal evolution shown in panel d).
@@ 6 Conclusions
In this work, data from the 2024 major flare campaign of Solar Orbiter were analysed. Those data showed that during the impulsive phase of the M2.5 GOES-class flare, SOL2024-03-23T23:41, when electron acceleration was strong, a significant fraction (approximately 50%) of the EUV kernels measured by HRIin the 174 Å passband were unresolved at a pixel scale of 135 km/pix and 2-pixel FWHM of the PSF core. These results are consistent with recent observations of chromospheric blobs in flare ribbons at high spatial resolution with SST and the DKIST (Thoen Faber et al., 2025; Yadav et al., 2025). In those works, chromospheric blobs with FWHM sizes in the range 140–200 km and 320–455 km were identified by SST and the DKIST, respectively. This corresponds to between 1–4 HRIpixels at 0.38 AU. Structures of this size are not resolvable by HRIEUV. Our results therefore agree with those obtained by SST and the DKIST, depsite the fact that HRIprobes a different layer of the solar atmosphere than the SST and the DKIST. In addition, we showed that for the flare studied, energy deposition occurred in a new region along the flare arcade in quick succession (on timescales of several seconds). Furthermore, an average EUV kernel time profile was derived. An upper bound was placed on the average half-maximum heating time of only a couple of seconds in duration, with correspondingly fast cooling times. Our best fit of the average EUV kernel time profile gave and . Together, these results highlight the transient and fragmented nature of solar flare energy release on small spatial scales. The findings presented here have important implications on our understanding of flare energy deposition. The full extent to which these observations alter our understanding warrants detailed modelling investigation. However, what remains strikingly clear is that the energy release process in flares is highly fragmented in space and time. Finally, we conclude that the bulk energy release process of solar flares consists of many smaller scale () events that behave like multiple flare-like energy release episodes occurring in quick succession.
@@ Acknowledgements.
We thank the anonymous referee for their constructive feedback and insightful comments. We are also grateful to Tom Van Doorsselaere for his contribution to the interpretation of the results presented here. Solar Orbiter is a space mission of international collaboration between ESA and NASA, operated by ESA. The EUI instrument was built by CSL, IAS, MPS, MSSL/UCL, PMOD/WRC, ROB, LCF/IO with funding from the Belgian Federal Science Policy Office (BELSPO/PRODEX PEA 4000112292 and 4000134088); the Centre National d’Etudes Spatiales (CNES); the UK Space Agency (UKSA); the Bundesministerium für Wirtschaft und Energie (BMWi) through the Deutsches Zentrum für Luft- und Raumfahrt (DLR); and the Swiss Space Office (SSO). HC and LAH had additional support from an EUI guest investigator grant issued by the EUI instrument team. The STIX instrument is an international collaboration between Switzerland, Poland, France, Czech Republic, Germany, Austria, Ireland, and Italy. LAH is supported by a Royal Society-Research Ireland University Research Fellowship (URFR1241775).References
@@
Appendix A A point source observed by an instrument with a PSF of 2-pixels at FWHM
In Figure 8 we demonstrate the effect of a 2-pixel PSF on a point source located at two extreme positions — at the centre of a pixel and at the intersection of four pixels (top and bottom row, respectively). The effect of applying a threshold is demonstrated in the final four columns. Naturally, as the threshold increases the number of pixels with non-zero counts reduces, decreasing the size of the kernel that would be identified from processed data. In this work, a fixed cut-off threshold is used for the entire time series. As a result, fainter kernels are more affected by artificial shrinkage. We therefore caution the reader not to over-interpret the shape of the size distributions shown in Figure 5 in the low-intensity range. This highlights the fact that point sources that have been artificially shrunk due to data preprocessing may appear as kernels with up to 9 pixels. Further, kernels with between 10 and 20 pixels are likely unresolved by HRIEUV.
Appendix B Kernel size histograms at cutoff
Figure 9 shows the frequency distribution of kernel sizes defined as the number of pixels within a kernel that have . Kernels whose watershed label did not extend beyond the 50% were flagged, in order to remove any bias introduced by using a fixed threshold. The distributions now have turnovers below 6 pixels because faint kernels whose boundaries were clipped by the fixed threshold have been filtered out of the dataset. Approximately 40% of kernels contain less than 6 pixels above the 50% peak kernel intensity level when the 5% threshold was applied to the data prior to segmentation on the running difference sequence. For comparison, a point source smeared out by an instrumental PSF with a 2-pixel FWHM contains 5 pixels above 50% of the peak kernel intensity when it falls on the centre of a pixel (see panel e) of Figure 8).
Appendix C Average EUV kernel time profiles